Consider the following reactions:

          3 Ni 2+(aq)  + 2Cr(OH)3 (s) + 10 OH –(aq)    →   3 Ni (s) + 2 CrO4 –2 (aq) + 8 H2O (l)

A) What is the value for n for this reactions?

B) Calculate the standard Gibbs Free-Energy

C) How long does it take a current of 5 A to deposit  95 g of nickel?



1
Expert's answer
2021-10-12T02:07:33-0400

A) Value for n for this reaction is 6


B) "E^0_{cell}=E^0_{Ni^{2+}}\/Ni+E^0_{Cr(OH)_3}\/CrO_4^{2-}=-0.257v+0.130v=-0.127v"

"\Delta G^o =-nFE^o=-6\times 96485\times (-0.127)=7.352157\times 10^4J"

So, the standard free energy is "73.5kJ"


C) Equivalent weight of Ni=molar mass"\/" valency=58.6934"\/2=29.3467g\/Eq"

Amount of charge required to deposit 95g of nickel"=96485C/1Eq\times 1Eq/29.3467g\times 95g=312337.503C"

time required, "t=q\/I=312337.503C\/5A=62467.5006s"

convert into hours "62467.5006\/3600=17.4hours"